2. End to or nestation af de mitis - Bad 2. Find the QR factorization of the matrix A = [1 1 2] 2 0 1 [2 1 7]
Answer #1

③ we want to factorization of find QR 7 = [a, a go Lora, a | 1 = 1 - = 1 - 4 - 5 - 3 1-2155 Y Fiz = 1= f* %7-1:1 ={(+)03+(3]©